OpenStudy (experimentx):

Find the value of \[ 1 + {1\cdot2\over1\cdot3} + {1\cdot2\cdot3\over1\cdot3\cdot5}+ {1\cdot2\cdot3\cdot4\over 1\cdot3\cdot5\cdot7} +\cdots \]

OpenStudy (anonymous):

*

OpenStudy (ganpat):

its like, n/n + (n)(n+1)/n(n+2) + (n)(n+1)(n+2)/n(n+2)(n+4) ... so a series numerator = n (n+1)/2.. addition of consecutive integers and denominator = n(n+1)... addition of all consecutive even integers so = n(n+1)/ 2(n)(n+1) So = 1/2... Not sure, just a try...

OpenStudy (experimentx):

\[ 1 + {\pi \over 2}\]

OpenStudy (anonymous):